OpenStudy (anonymous):

If t1 = 4 and tn = tn-1 + negative five-halvesfind the first 4 terms of the sequence

jimthompson5910 (jim_thompson5910):

t1 is the first term saying t1 = 4 means that the first term is 4

jimthompson5910 (jim_thompson5910):

tn or \(\large t_{n}\) is the nth term t(n-1) or \(\large t_{n-1}\) is the term just before the nth term

jimthompson5910 (jim_thompson5910):

so what this equation means \[\Large t_{n} = t_{n-1} + \frac{-5}{2}\] is that you add on -5/2 to any term to get the next term

jimthompson5910 (jim_thompson5910):

how do we find t2 (the second term)? we add -5/2 to the first term to get 4 + (-5/2) = ??

OpenStudy (anonymous):

the answers are a. 4, -1/2, -4, 9/2 b. 4, 1/2, -3, -7/2 c. 4, 1/2, -2, -7/2 d. 4, 3/2, -1, -7/2

jimthompson5910 (jim_thompson5910):

what is 4 + (-5/2) equal to

OpenStudy (anonymous):

6.5?

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

the 5/2 is negative, be careful not to lose the negative sign

jimthompson5910 (jim_thompson5910):

4 + (-5/2) is the same as saying 4 - (5/2)

OpenStudy (anonymous):

oh okay but how do ifind the answer im not sure

jimthompson5910 (jim_thompson5910):

what is 4 - (5/2) equal to

jimthompson5910 (jim_thompson5910):

you can think of 5/2 as 2.5 if you want so 4 - (5/2) = 4 - 2.5 = ??

OpenStudy (anonymous):

1.5

jimthompson5910 (jim_thompson5910):

sorry my notifications are all messed up and I'm just noticing this message

jimthompson5910 (jim_thompson5910):

anyways, yes 4 - 2.5 = 1.5 convert 1.5 to a fraction to get 1.5 = 3/2

jimthompson5910 (jim_thompson5910):

first term is 4 second term is 3/2

jimthompson5910 (jim_thompson5910):

to get the third term, you subtract 5/2 from 3/2 3/2 - 5/2 = ???